# Towards a robust frequency-domain analysis: Spectral Rényi divergence revisited

### v1

@@ Appendix A Proof of Proposition 2
Appendix B Proof of Theorem 3
Proof.
We prove Theorem 3 by the mathematical induction.
Step 1: Consider $m=1$. Observe
$=\frac{\gamma_{1}\alpha}{(1-\alpha)n}\left(\frac{S_{\theta^{(0)}}(\omega^{*})}{\alpha S_{\theta^{(0)}}(\omega^{*})+(1-\alpha)\widetilde{I}_{n}(\omega^{*})}-\frac{S_{\theta^{(0)}}(\omega^{*})}{\alpha S_{\theta^{(0)}}(\omega^{*})+(1-\alpha)(\widetilde{I}_{n}(\omega^{*})+z)}\right)\nabla_{\theta^{(0)}}\log S_{\theta}(\omega^{*})$
This yields
$\left\|\widehat{\theta}^{(1)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]-\widehat{\theta}^{(1)}_{\alpha}[\widetilde{I}_{n}]\right\|\leq\frac{1}{n}\frac{\gamma_{1}}{1-\alpha}\|\nabla_{\theta^{(0)}}S_{\theta}(\omega^{*})\|,$
which proves the assertion for $m=1$.
By the cancelling technique and by the triangle inequality, we get
=$=\left\|\left\{\widehat{\theta}^{(m+1)}_{\alpha}[\widetilde{I}_{n}^{z,\omega^{*}}]-\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}^{z,\omega^{*}}]\right\}-\left\{\widehat{\theta}^{(m+1)}_{\alpha}[\widetilde{I}_{n}]-\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]\right\}-\left\{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}^{z,\omega^{*}}]-\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]\right\}\right\|$
$\leq\left\|\mathcal{G}_{\alpha}\left(\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]\,;\,\widetilde{I}^{z,\omega^{*}}_{n}\right)-\mathcal{G}_{\alpha}\left(\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]\,;\,\widetilde{I}_{n}\right)\right\|+\varepsilon.$
$=\frac{\gamma_{m}\alpha}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\left\{1-\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}\right\}\nabla_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}\log S_{\theta}(\omega)$
$=\frac{\gamma_{m}\alpha}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\left\{1-\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}\right\}\nabla_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}\log S_{\theta}(\omega)$
$\quad+\frac{\gamma_{m}\alpha}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\left\{1-\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}\right\}\nabla_{t}^{2}\log S_{\theta}(\omega)(\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]-\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}])$
$=\frac{\gamma_{m}\alpha}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\left\{1-\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}\right\}\nabla_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}\log S_{\theta}(\omega)$
$\quad+O\left(\frac{2\gamma_{m}\varepsilon}{(1-\alpha)n}\sup_{\theta\in\Theta}\sum_{\omega\in\Omega_{n}}\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\right),$
$=\frac{\gamma_{m}\alpha}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\left\{\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)+(1-\alpha)\widetilde{I}_{n}(\omega)}-\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}\right\}$
$\qquad\qquad\quad\qquad\qquad\nabla_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}\log S_{\theta}(\omega)$
$\quad+O\left(\frac{2\gamma_{m}\varepsilon}{(1-\alpha)n}\sup_{\theta\in\Theta}\sum_{\omega\in\Omega_{n}}\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\right)$
$=\frac{\gamma_{m}\alpha}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\left\{\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)+(1-\alpha)\widetilde{I}_{n}(\omega)}-\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}\right.$
$\left.\qquad\qquad\quad\qquad\qquad+\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}-\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}\right\}$
$\qquad\qquad\quad\qquad\qquad\nabla_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}\log S_{\theta}(\omega)$
$\quad+O\left(\frac{2\gamma_{m}\varepsilon}{(1-\alpha)n}\sup_{\theta\in\Theta}\sum_{\omega\in\Omega_{n}}\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\right)$
$=\frac{\gamma_{m}\alpha}{(1-\alpha)n}(T_{1}+T_{2})+O\left(\frac{2\gamma_{m}\varepsilon}{(1-\alpha)n}\sup_{\theta\in\Theta}\sum_{\omega\in\Omega_{n}}\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\right),$
where
$T_{1}$ $:=\sum_{\omega\in\Omega_{n}}\left\{\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)+(1-\alpha)\widetilde{I}_{n}(\omega)}-\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}\right\}\nabla_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}\log S_{\theta}(\omega),$
$T_{2}$ $:=\sum_{\omega\in\Omega_{n}}\left\{\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}-\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}\right\}$
$\qquad\qquad\quad\nabla_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}\log S_{\theta}(\omega).$
Here we have
$T_{1}$ $=\frac{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega^{*})}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega^{*})+(1-\alpha)\widetilde{I}_{n}(\omega^{*})}\frac{(1-\alpha)z}{(1-\alpha)z+\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega^{*})+(1-\alpha)\widetilde{I}_{n}(\omega^{*})}\frac{\nabla_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}\log S_{\theta}(\omega^{*})}{\alpha}$
and
$T_{2}$ $=\sum_{\omega\in\Omega_{n}}\frac{(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)-S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}\nabla_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}\log S_{\theta}(\omega)$
Then we obtain
$\quad+\varepsilon\left\{\frac{\gamma_{m}}{(1-\alpha)}\left(\sup_{\theta,\omega}\|\nabla_{\theta,\omega}S_{\theta}(\omega)\|\right)^{2}+\frac{2\gamma_{m}}{(1-\alpha)}\sup_{\theta\in\Theta}\frac{1}{n}\sum_{\omega\in\Omega_{n}}\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\right\}$
and thus
$\quad+\varepsilon\left\{1+\frac{\gamma_{m}}{(1-\alpha)}\left(\sup_{\theta,\omega}\|\nabla_{\theta,\omega}S_{\theta}(\omega)\|\right)^{2}+\frac{2\gamma_{m}}{(1-\alpha)}\sup_{\theta\in\Theta}\frac{1}{n}\sum_{\omega\in\Omega_{n}}\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\right\}.$
Appendix C Additional simulation studies
@@
Table 3: The mean values of biases with standard deviations without any trend. For each initial value, the values closest to zero are underlined.
Initial value $\widehat{\sigma}-\sigma^{*}$ $\widehat{\omega_{c}}-\omega_{c}^{*}$ $\widehat{Q}-Q^{*}$
Rényi ($\alpha=0.50$) $\theta^{(0),4}$0.03 ($\pm$0.05) -0.11 ($\pm$0.07) 0.26 ($\pm$0.15)
Rényi ($\alpha=0.75$) $\theta^{(0),4}$ 0.04 ($\pm$0.06) -0.30 ($\pm$0.03) 1.33 ($\pm$0.17)
Rényi ($\alpha=0.90$) $\theta^{(0),4}$ 0.09 ($\pm$0.08) -0.39 ($\pm$0.03) 6.13 ($\pm$0.04)
Itakura–Saito with $I_{n}$ $\theta^{(0),4}$ $5\cdot 10^{15}$ (±$\pm 8\cdot 10^{14}$) $1\cdot 10^{16}$ (±$\pm 1\cdot 10^{15}$) $8\cdot 10^{16}$ (±$\pm 1\cdot 10^{16}$)
Itakura–Saito with $I_{n}^{S}$ $\theta^{(0),4}$ $5\cdot 10^{18}$ (±$\pm 9\cdot 10^{18}$) $1\cdot 10^{19}$ (±$\pm 1\cdot 10^{19}$) $9\cdot 10^{19}$ (±$\pm 1\cdot 10^{20}$)
Rényi ($\alpha=0.50$) $\theta^{(0),5}$0.01($\pm$0.06) -0.18 ($\pm$0.05) 0.47 ($\pm$0.16)
Rényi ($\alpha=0.75$) $\theta^{(0),5}$ 0.01 ($\pm$ 0.06) -0.21 ($\pm$0.05) 0.68 ($\pm$0.22)
Rényi ($\alpha=0.90$) $\theta^{(0),5}$ 0.02 ($\pm$0.06) -0.24 ($\pm$0.04) 1.18 ($\pm$0.34)
Itakura–Saito with $I_{n}$ $\theta^{(0),5}$ 0.07 ($\pm$0.11) -0.08 ($\pm$0.12) 0.12 ($\pm$0.18)
Itakura–Saito with $I_{n}^{S}$ $\theta^{(0),5}$0.28 ($\pm$0.16) 1.84 ($\pm$1.66) 1.90 ($\pm$0.63)
Table 4: The mean values of biases with standard deviations with the trigonometric trends. For each initial value, the values closest to zero are underlined.
Initial value $\widehat{\sigma}-\sigma^{*}$ $\widehat{\omega_{c}}-\omega_{c}^{*}$ $\widehat{Q}-Q^{*}$
Rényi ($\alpha=0.50$) $\theta^{(0),4}$ 0.21 ($\pm$0.07) -0.21 ($\pm$0.06) 0.24 ($\pm$0.14)
Rényi ($\alpha=0.75$) $\theta^{(0),4}$ 0.61 ($\pm$0.09) -0.43 ($\pm$0.02) 1.14 ($\pm$0.20)
Rényi ($\alpha=0.90$) $\theta^{(0),4}$ 1.65 ($\pm$0.12) -0.60 ($\pm$0.02) 6.05 ($\pm$0.05)
Itakura–Saito with $I_{n}$ $\theta^{(0),4}$ $5\cdot 10^{15}$ (±$\pm 8\cdot 10^{14}$) $1\cdot 10^{16}$ (±$\pm 1\cdot 10^{15}$) $8\cdot 10^{16}$ (±$\pm 1\cdot 10^{16}$)
Itakura–Saito with $I_{n}^{S}$ $\theta^{(0),4}$ $5\cdot 10^{18}$ (±$\pm 9\cdot 10^{18}$) $1\cdot 10^{19}$ (±$\pm 1\cdot 10^{19}$) $9\cdot 10^{19}$ (±$\pm 1\cdot 10^{20}$)
Rényi ($\alpha=0.50$) $\theta^{(0),5}$ 0.23 ($\pm$0.07) -0.26 ($\pm$0.04) 0.41 ($\pm$0.16)
Rényi ($\alpha=0.75$) $\theta^{(0),5}$ 0.55 ($\pm$0.08) -0.35 ($\pm$0.04) 0.55 ($\pm$0.20)
Rényi ($\alpha=0.90$) $\theta^{(0),5}$ 1.50 ($\pm$0.10) -0.50 ($\pm$0.03) 0.85 ($\pm$0.35)
Itakura–Saito with $I_{n}$ $\theta^{(0),5}$ 3.91 ($\pm$0.07) -0.34 ($\pm$0.02) -0.40 ($\pm$0.02)
Itakura–Saito with $I_{n}^{S}$ $\theta^{(0),5}$ 2.85 ($\pm$0.95) -0.17 ($\pm$0.51) 2.26 ($\pm$0.93)
Figures 3 and 4 display an instance of a set of the estimated spectral densities. In each figure, the following representations are used:

### v3

@@ Appendix D Proof of Proposition 3
Appendix E Additional simulation studies
@@
Table 3: The mean values of biases with standard deviations without any trend. For each initial value, the values closest to zero are underlined. Rényi is abbreviated as R; Itakura–Saito is abbreviated as IS; Initial value is abbreviated as Init. Values greater than 10 are denoted by *.
Init $\widehat{\sigma}-\sigma^{*}$ $\widehat{\omega_{c}}-\omega_{c}^{*}$ $\widehat{Q}-Q^{*}$
R($\alpha=0.50$) $\theta^{(0),4}$0.03 ($\pm$0.05) -0.11 ($\pm$0.07) 0.26 ($\pm$0.15)
R ($\alpha=0.75$) $\theta^{(0),4}$ 0.04 ($\pm$0.06) -0.30 ($\pm$0.03) 1.33 ($\pm$0.17)
R ($\alpha=0.90$) $\theta^{(0),4}$ 0.09 ($\pm$0.08) -0.39 ($\pm$0.03) 6.13 ($\pm$0.04)
IS with $I_{n}$ $\theta^{(0),4}$
IS with $I_{n}^{S}$ $\theta^{(0),4}$
R ($\alpha=0.50$) $\theta^{(0),5}$0.01($\pm$0.06) -0.18 ($\pm$0.05) 0.47 ($\pm$0.16)
R ($\alpha=0.75$) $\theta^{(0),5}$ 0.01 ($\pm$ 0.06) -0.21 ($\pm$0.05) 0.68 ($\pm$0.22)
R ($\alpha=0.90$) $\theta^{(0),5}$ 0.02 ($\pm$0.06) -0.24 ($\pm$0.04) 1.18 ($\pm$0.34)
IS with $I_{n}$ $\theta^{(0),5}$ 0.07 ($\pm$0.11) -0.08 ($\pm$0.12) 0.12 ($\pm$0.18)
IS with $I_{n}^{S}$ $\theta^{(0),5}$0.28 ($\pm$0.16) 1.84 ($\pm$1.66) 1.90 ($\pm$0.63)
Table 4: The mean values of biases with standard deviations with the trigonometric trends. For each initial value, the values closest to zero are underlined. Rényi is abbreviated as R; Itakura–Saito is abbreviated as IS; Initial value is abbreviated as Init.
Init $\widehat{\sigma}-\sigma^{*}$ $\widehat{\omega_{c}}-\omega_{c}^{*}$ $\widehat{Q}-Q^{*}$
R ($\alpha=0.50$) $\theta^{(0),4}$ 0.21 ($\pm$0.07) -0.21 ($\pm$0.06) 0.24 ($\pm$0.14)
R ($\alpha=0.75$) $\theta^{(0),4}$ 0.61 ($\pm$0.09) -0.43 ($\pm$0.02) 1.14 ($\pm$0.20)
R ($\alpha=0.90$) $\theta^{(0),4}$ 1.65 ($\pm$0.12) -0.60 ($\pm$0.02) 6.05 ($\pm$0.05)
IS with $I_{n}$ $\theta^{(0),4}$
IS with $I_{n}^{S}$ $\theta^{(0),4}$
R ($\alpha=0.50$) $\theta^{(0),5}$ 0.23 ($\pm$0.07) -0.26 ($\pm$0.04) 0.41 ($\pm$0.16)
R ($\alpha=0.75$) $\theta^{(0),5}$ 0.55 ($\pm$0.08) -0.35 ($\pm$0.04) 0.55 ($\pm$0.20)
R ($\alpha=0.90$) $\theta^{(0),5}$ 1.50 ($\pm$0.10) -0.50 ($\pm$0.03) 0.85 ($\pm$0.35)
IS with $I_{n}$ $\theta^{(0),5}$ 3.91 ($\pm$0.07) -0.34 ($\pm$0.02) -0.40 ($\pm$0.02)
IS with $I_{n}^{S}$ $\theta^{(0),5}$ 2.85 ($\pm$0.95) -0.17 ($\pm$0.51) 2.26 ($\pm$0.93)
Figures 4 and 5 display an instance of a set of the estimated spectral densities. In each figure, the following representations are used: